# Temporal Valued Constraint Satisfaction Problems

### v1

@@ 5. Temporal VCSPs
5.1. Preliminaries on temporal CSPs
We first define several important relations on $\mathbb{Q}$ that already played a role in the classification of temporal CSPs [9].
Definition 5.1.
@@ 5.1. Preliminaries on temporal CSPs
$\operatorname{Betw}$ :=$:=\{(x,y,z)\in\mathbb{Q}^{3}\mid(x<y<z)\vee(z<y<x)\},$
$\operatorname{Cycl}$ :=$:=\{(x,y,z)\in\mathbb{Q}^{3}\mid(x<y<z)\vee(y<z<x)\vee(z<x<y)\},$
$\operatorname{Sep}$ :=$:=\{(x_{1},y_{1},x_{2},y_{2})\in\mathbb{Q}^{4}\mid\begin{aligned}&(x_{1}<x_{2}<y_{1}<y_{2})\vee(x_{1}<y_{2}<y_{1}<x_{2})\vee(y_{1}<x_{2}<x_{1}<y_{2})\\ \vee~{}&(y_{1}<y_{2}<x_{1}<x_{2})\vee(x_{2}<x_{1}<y_{2}<y_{1})\vee(x_{2}<y_{1}<y_{2}<x_{1})\\ \vee~{}&(y_{2}<x_{1}<x_{2}<y_{1})\vee(y_{2}<y_{1}<x_{2}<x_{1})\},\end{aligned}$
$T_{3}$ :=$:=\{(x,y,z)\in\mathbb{Q}^{3}\mid(x=y<z)\vee(x=z<y)\}.$
Theorem 5.2 (Theorem 20 in [9]).
@@
  • There is a primitive positive definition of $<$ in ${\mathfrak{A}}$.
We need the following operations on $\mathbb{Q}$. By $\min$ and $\max$ we refer to the binary minimum and maximum operation on the set $\mathbb{Q}$, respectively.
Definition 5.3.
Let $e_{<0},e_{>0}$ be any endomorphisms of $(\mathbb{Q};<)$ satisfying $e_{<0}(x)<0$ and $e_{>0}(x)>0$ for every $x\in\mathbb{Q}$. We denote by $\pi\pi$ the binary operation on $\mathbb{Q}$ defined by
$\pi\pi(x,y)=\begin{cases}e_{<0}(x)&x\leq 0,\\ e_{>0}(y)&x>0.\end{cases}$
The operation $\operatorname{lex}$ is any binary operation on $\mathbb{Q}$ satisfying $\operatorname{lex}(x,y)<\operatorname{lex}(x^{\prime},y^{\prime})$ iff $x<x^{\prime}$, or $x=x^{\prime}$ and $y<y^{\prime}$ for all $x,x^{\prime},y,y^{\prime}\in\mathbb{Q}$. We denote by $\operatorname{ll}$ the binary operation on $\mathbb{Q}$ defined by
$\operatorname{ll}(x,y)=\begin{cases}\operatorname{lex}(e_{<0}(x),e_{<0}(y))&x\leq 0,\\ \operatorname{lex}(e_{>0}(y),e_{>0}(x))&x>0.\end{cases}$
Definition 5.4.
Let $e_{<}$, $e_{=}$ and $e_{>}$ be any endomorphisms of $(\mathbb{Q};<)$ satisfying for all $x,\varepsilon\in\mathbb{Q}$, $\varepsilon>0$,
$e_{=}(x)<e_{>}(x)<e_{<}(x)<e_{=}(x+\varepsilon).$
We denote by $\operatorname{mi}$ the binary operation on $\mathbb{Q}$ defined by
$\operatorname{mi}(x,y)=\begin{cases}e_{<}(x)&x<y,\\ e_{=}(x)&x=y,\\ e_{>}(y)&x>y.\end{cases}$
Definition 5.5.
Let $e_{=}$ and $e_{\neq}$ be any endomorphisms of $(\mathbb{Q};<)$ satisfying for all $x,\varepsilon\in\mathbb{Q}$, $\varepsilon>0$,
$e_{\neq}(x)<e_{=}(x)<e_{\neq}(x+\varepsilon).$
We denote by $\operatorname{mx}$ the binary operation on $\mathbb{Q}$ defined by
$\operatorname{mx}(x,y)=\begin{cases}e_{\neq}(\min(x,y))&x\neq y,\\ e_{=}(x)&x=y.\end{cases}$
The construction of endomorphisms that appear in Definitions 5.4 and 5.5 can be found for example in [3, Section 12.5]. The following was observed and used in [9].
Lemma 5.6.
@@
Definition 5.7.
@@
$-R=\{(-a_{1},\dots,-a_{\ell})\mid(a_{1},\dots,a_{\ell})\in R\}.$
Note that $\min^{*}=\max$ and the relation (>)-(>)- ( > ) is equal to $<$. Statements about operations and relations on $\mathbb{Q}$ can be naturally dualized and we will often refer to the dual version of a statement.
By combining Theorem 50, Corollary 51, Corollary 52 and the accompanying remarks in [9], we obtain the following; see also Theorem 12.10.1 in [3].
Theorem 5.8.
Let ${\mathfrak{A}}$ be a relational structure such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Then exactly one of the following is true.
  1. (1)
@@
  2. (2)
    ${\mathfrak{A}}$ primitively positively defines one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $-T_{3}$, or $\operatorname{Dis}$. In this case, ${\mathfrak{A}}$ has a reduct ${\mathfrak{A}}^{\prime}$ with a finite signature such that $\operatorname{CSP}({\mathfrak{A}}^{\prime})$ is NP-complete.
Moreover, it is decidable whether (1) or (2) holds.
@@
Theorem 5.9 (Theorem 12.0.1 in [3]; see also Theorem 7.24 in [14]).
@@
    $\operatorname{Pol}({\mathfrak{A}})$ pp-constructs $K_{3}$. In this case, there exists a reduct ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$ with a finite signature such that $\operatorname{CSP}({\mathfrak{A}}^{\prime})$ is NP-complete.
Proposition 5.10.
Each of the relational structures $(\mathbb{Q};\operatorname{Betw}),(\mathbb{Q};\operatorname{Cycl}),(\mathbb{Q};\operatorname{Sep}),(\mathbb{Q};T_{3})$, and $(\mathbb{Q};-T_{3})$ pp-constructs $K_{3}$.
Proof.
@@
This is proved in the proof of [3, Theorem 12.0.1]. In fact, the proof shows that each of these structures pp-interprets all finite structures. Since $K_{3}$ is finite and a pp-interpretation is a special case of a pp-construction, the statement follows. ∎
Proposition 5.11 (Proposition 25, 27, and 29 in [9]).
Proposition 5.12 (Lemma 12.4.4 in [3]).
Let ${\mathfrak{A}}$ be a relational structure such that $\operatorname{Pol}({\mathfrak{A}})$ contains $\operatorname{lex}$ and $\pi\pi$. Then $\operatorname{Pol}({\mathfrak{A}})$ contains $\operatorname{ll}$.
Note that there exists $\alpha\in\operatorname{Aut}(\mathbb{Q};<)$ such that, for all $x,y\in\mathbb{Q}$, $\operatorname{lex}^{*}(x,y)=\alpha(\operatorname{lex}(x,y))$. Hence, whenever ${\mathfrak{A}}$ is a relational structure such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$, then $\operatorname{Pol}({\mathfrak{A}})$ contains $\operatorname{lex}$ if and only if it contains $\operatorname{lex}^{*}$. This is relevant for dualising statements like Proposition 5.12. We also need the following result from [5].
Theorem 5.13 (Theorem 5.1 in [5]).
@@
Let ${\mathfrak{A}}$ be a relational structure such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. If $\pi\pi\in\operatorname{Pol}({\mathfrak{A}})$ and $\operatorname{ll}\notin\operatorname{Pol}({\mathfrak{A}})$, then the relation
has a primitive positive definition in ${\mathfrak{A}}$.
5.2. Expressibility of Valued Relations
In this section we consider valued structures ${\mathfrak{A}}$ such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$, and study expressibility of valued relations in ${\mathfrak{A}}$. For $\alpha,\beta,\gamma\in{\mathbb{Q}}\cup\{\infty\}$, define the binary valued relation $R_{\alpha,\beta,\gamma}$ on ${\mathbb{Q}}$:

### v2

@@ 5 Temporal VCSPs
5.1 Preliminaries on temporal CSPs
We first define several important relations on $\mathbb{Q}$ that already played a role in the classification of temporal CSPs [8]; in this paper, we treat these relations in a black-box fashion.
Definition 34
@@ 5.1 Preliminaries on temporal CSPs
$\operatorname{Betw}$ $:=\{(x,y,z)\in\mathbb{Q}^{3}\mid(x<y<z)\vee(z<y<x)\},$
$\operatorname{Cycl}$ $:=\{(x,y,z)\in\mathbb{Q}^{3}\mid(x<y<z)\vee(y<z<x)\vee(z<x<y)\},$
$\operatorname{Sep}$ $:=\{(x_{1},y_{1},x_{2},y_{2})\in\mathbb{Q}^{4}\begin{aligned} \mid~&(x_{1}<x_{2}<y_{1}<y_{2})\vee(x_{1}<y_{2}<y_{1}<x_{2})\\ \vee~&(y_{1}<x_{2}<x_{1}<y_{2})\vee(y_{1}<y_{2}<x_{1}<x_{2})\\ \vee~&(x_{2}<x_{1}<y_{2}<y_{1})\vee(x_{2}<y_{1}<y_{2}<x_{1})\\ \vee~&(y_{2}<x_{1}<x_{2}<y_{1})\vee(y_{2}<y_{1}<x_{2}<x_{1})\},\end{aligned}$
$T_{3}$ $:=\{(x,y,z)\in\mathbb{Q}^{3}\mid(x=y<z)\vee(x=z<y)\}.$
Theorem 35 (Theorem 20 in [8])
@@
  • 4.
    There is a primitive positive definition of << in $\mathfrak{A}$}.
To build on the results on temporal CSPs, we need the following operations on $\mathbb{Q}$; they will be used in a black-box fashion as well. By $\min$ and $\max$ we refer to the binary minimum and maximum operation on the set $\mathbb{Q}$, respectively.
Definition 36
Let $e_{<0},e_{>0}$ be endomorphisms of $(\mathbb{Q};<)$ satisfying $e_{<0}(x)<0$ and $e_{>0}(x)>0$ for every $x\in\mathbb{Q}$. We denote by $\pi\pi$ the binary operation on $\mathbb{Q}$ defined by
$\pi\pi(x,y)=\begin{cases}e_{<0}(x)&\text{if }x\leq 0,\\ e_{>0}(y)&\text{if }x>0.\end{cases}$
Definition 37
Let $e_{<0},e_{>0}$ be endomorphisms of $(\mathbb{Q};<)$ satisfying $e_{<0}(x)<0$ and $e_{>0}(x)>0$ for every $x\in\mathbb{Q}$. The operation $\operatorname{lex}$ is any binary operation on $\mathbb{Q}$ satisfying $\operatorname{lex}(x,y)<\operatorname{lex}(x^{\prime},y^{\prime})$ iff $x<x^{\prime}$, or $x=x^{\prime}$ and $y<y^{\prime}$ for all $x,x^{\prime},y,y^{\prime}\in\mathbb{Q}$. We denote by $\operatorname{ll}$ the operation on $\mathbb{Q}$ defined by
$\operatorname{ll}(x,y)=\begin{cases}\operatorname{lex}(e_{<0}(x),e_{<0}(y))&\text{if }x\leq 0,\\ \operatorname{lex}(e_{>0}(y),e_{>0}(x))&\text{if }x>0.\end{cases}$
Definition 38
Let $_{<}, e=e_{=}$ and $e_{>}$ be endomorphisms of $(\mathbb{Q};<)$ satisfying for all $x,\varepsilon\in\mathbb{Q}$, $\varepsilon>0$, $e_{=}(x)<e_{>}(x)<e_{<}(x)<e_{=}(x+\varepsilon).$ We denote by $\operatorname{mi}$ the binary operation on $\mathbb{Q}$ defined by
$\operatorname{mi}(x,y)=\begin{cases}e_{<}(x)&\text{if }x<y,\\ e_{=}(x)&\text{if }x=y,\\ e_{>}(y)&\text{if }x>y.\end{cases}$
Definition 39
Let $e_{=}$ and $e_{\neq}$ be any endomorphisms of $(\mathbb{Q};<)$ satisfying for all $x,\varepsilon\in\mathbb{Q}$, $\varepsilon>0$, $e_{\neq}(x)<e_{=}(x)<e_{\neq}(x+\varepsilon).$ We denote by $\operatorname{mx}$ the binary operation on $\mathbb{Q}$ defined by
$\operatorname{mx}(x,y)=\begin{cases}e_{\neq}(\min(x,y))&\text{if }x\neq y,\\ e_{=}(x)&\text{if }x=y.\end{cases}$
The construction of endomorphisms that appear in Definitions 38 and 39 can be found for example in [17, Section 12.5]. We now give a few technical results that will be useful for the proofs in the upcoming sections. The following lemma was observed and used in [8].
Lemma 40
@@
Proposition 41 (Proposition 25, 27, and 29 in [8])
Theorem 42 ([4, Theorem 5.1])
Let $\mathfrak{A}$} be a relational structure with $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Aut}(\mathbb{Q};<)$ and suppose that $\mathfrak{A}$} contains the relation <<. If $\pi\pi\in\operatorname{Pol}({\mathfrak{A}})$ and $\operatorname{ll}\notin\operatorname{Pol}({\mathfrak{A}})$, then the relation
has a primitive positive definition in $\mathfrak{A}$}.
Definition 43
@@
$R=\{(-a_{1},\dots,-a_{\ell})\mid(a_{1},\dots,a_{\ell})\in R\}.$
Note that $\min^{*}=\max$ and the relation (>)-(>) is equal to <<. Statements about operations and relations on $\mathbb{Q}$ can be naturally dualised and we may refer to the dual version of a statement.
By combining results from [8] (Theorem 50, Corollary 51, Corollary 52, and the accompanying remarks), we obtain the following; also see Theorem 12.10.1 in [17].
Theorem 44
Let $\mathfrak{A}$} be a relational structure with $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Then exactly one of the following is true.
  1. 1.
@@
  2. 2.
    $\mathfrak{A}$} primitively positively defines one of the relations $\operatorname{Betw}$, $\operatorname{Cycl}$, $\operatorname{Sep}$, $T_{3}$, $T_{3}$, or $\operatorname{Dis}$. In this case, $\mathfrak{A}$} has a reduct $\mathfrak{A}$$^{\prime}$ with a finite signature such that $\operatorname{CSP}({\mathfrak{A}}^{\prime})$ is NP-complete.
Moreover, if $\mathfrak{A}$} has a finite signature, then it is decidable whether item 1 or item 2 holds.
@@
Theorem 45 ([17, Theorem 12.0.1]; see also [14, Theorem 7.24])
@@
    $\mathfrak{A}$} $p$-constructs $K_{3}$. In this case, there exists a reduct $\mathfrak{A}$$^{\prime}$ of $\mathfrak{A}$} with a finite signature such that $\operatorname{CSP}({\mathfrak{A}}^{\prime})$ is NP-complete.
The following proposition relates hardness of temporal CSPs to pp-constructing $K_{3}$; it is essentially proven in [17, Theorem 12.0.1].
Proposition 46
The relational structures $(\mathbb{Q};\operatorname{Betw})$, $(\mathbb{Q};\operatorname{Cycl})$, $(\mathbb{Q};\operatorname{Sep})$, $(\mathbb{Q};T_{3})$, and $(\mathbb{Q};-T_{3})$ all pp-construct $K_{3}$.
Proof 9
@@
This is proven in the proof of [17, Theorem 12.0.1]. In fact, the proof shows that each of these structures pp-interprets all finite structures. Since $K_{3}$ is finite and a pp-interpretation is a special case of a pp-construction, the statement follows. $\qed$
5.2 A polynomial-time tractability result
In this section we present the only non-trivial classes of polynomial-time tractable temporal VCSPs. In Lemma 47 below we present a polynomial-time algorithm for VCSPs of valued structures $\mathfrak{A}$} improved by $\operatorname{ll}$ or $\operatorname{ll}^{*}$.
Lemma 47
@@ 5.2 A polynomial-time tractability result
Let $\mathfrak{A}$} be a valued structure over a finite signature such that $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$. Suppose that $\operatorname{ll}\in\operatorname{fPol}({\mathfrak{A}})$ or $\operatorname{ll}^{*}\in\operatorname{fPol}({\mathfrak{A}})$. Then $\operatorname{VCSP}({\mathfrak{A}})$ is in P.
In Algorithm 1 we provide the polynomial-time algorithm that is described in the proof of Lemma 47. The high level idea of the algorithm is similar to the algorithm in Lemma 29: we detect forced equalities and collapse the corresponding variables, then we find an injective solution to the modified instance.
Input: An instance $(\phi,u)$ of $\operatorname{VCSP}({\mathfrak{A}})$ with variables $V=\{v_{1},\dots,v_{N}\}$
Output: true or false
// $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$ is solved by the algorithm from [8] based on $\operatorname{ll}$ or $\operatorname{ll}^{*}$
if $\phi$ satisfiable over $\operatorname{Feas}({\mathfrak{A}})$ then
[…]
else
 return false
$\phi^{\prime}:=\phi$ where $v_{j}$ is replaced by $v_{i}$ for every $(v_{i},v_{j})\in E$ with $i<j$
$\phi^{\prime}=\phi_{1}^{\prime}+\dots+\phi_{n}^{\prime}$ where $\phi_{j}^{\prime}$ are atomic
for $j\in[n]$ do
 $\phi_{j}^{\prime}=R(x_{1},\dots,x_{k})$ for some $R\in\tau$
   let $y_{1}^{j},\dots y_{\ell_{j}}^{j}$ be an enumeration of all distinct variables in $\{x_{1},\dots,x_{k}\}$
 $S_{j}(y_{1}^{j},\dots y_{\ell_{j}}^{j}):=R(x_{1},\dots x_{k})$
${\mathfrak{B}}:=(\mathbb{Q};\operatorname{Opt}(S_{1}),\dots,\operatorname{Opt}(S_{N}))$
// by the assumptions $\operatorname{CSP}({\mathfrak{B}})$ is in P
$\psi:=\operatorname{Opt}(S_{1})(y_{1}^{1},\dots y_{\ell_{1}}^{1})\wedge\dots\wedge\operatorname{Opt}(S_{n})(y_{1}^{n},\dots y_{\ell_{n}}^{n})$
// it is implied that $\psi$ is satisfiable over $\mathfrak{B}$}
$f:=$ solution to $\psi$ over $\mathfrak{B}$}
if the cost of $\phi^{\prime}$ evaluated under $f$ is $\leq u$ then
 return true
else
 return false
Algorithm 1 An algorithm for $\operatorname{VCSP}({\mathfrak{A}})$ where $\mathfrak{A}$} is improved by $\operatorname{ll}$ or $\operatorname{ll}^{*}$.
Proof 10 (Lemma 47)
Let $R\in\langle{\mathfrak{A}}\rangle$ be of arity $k$. Since $\operatorname{ll}$ or $\operatorname{ll}^{*}$ lies in $\operatorname{fPol}({\mathfrak{A}})$, $\operatorname{lex}\in\operatorname{fPol}({\mathfrak{A}})$ by Lemma 40, and, in particular, $\operatorname{lex}$ improves $R$ by Lemma 18. Therefore, for every injective tuple $s\in\mathbb{Q}^{k}$ and any $t\in\mathbb{Q}^{k}$, it holds that
$R(s)=R(\operatorname{lex}(s,t))\leq 1/2\cdot(R(s)+R(t)),$
Let $(\phi,u)$ be an instance of $\operatorname{VCSP}({\mathfrak{A}})$ with variable set $V=\{v_{1},\dots,v_{N}\}$. Note that if we interpret $\phi$ over $\operatorname{Feas}({\mathfrak{A}})$, we can view it as an instance of $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$ where each summand $R(x_{1},\dots,x_{k})$ of $\phi$ is interpreted as $\operatorname{Feas}(R^{\mathfrak{A}})(x_{1},\dots,x_{k})$. By the assumption and Lemma 18, $\operatorname{Feas}({\mathfrak{A}})$ is preserved by $\operatorname{ll}$ or $\operatorname{ll}^{*}$. Hence, by Theorem 44, $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$ is solvable in polynomial time and we can use the polynomial-time algorithm from [8] based on the operation $\operatorname{ll}$ or $\operatorname{ll}^{*}$ to solve $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$. If $\phi$, viewed as a primitive positive formula, is not satisfiable over $\operatorname{Feas}({\mathfrak{A}})$, then the minimum of $\phi$ is above every rational threshold and $(\phi,u)$ is rejected. Otherwise, we may compute the set $E\subseteq V^{2}$ of all pairs $(x,y)$ such that $f(x)=f(y)$ in every solution of $f\colon V\to\mathbb{Q}$ of $\phi$ over $\operatorname{Feas}({\mathfrak{A}})$ (we may assume without loss of generality that $\operatorname{Feas}({\mathfrak{A}})$ contains the relation $(\neq)_{0}^{\infty}$; since $\operatorname{Feas}({\mathfrak{A}})$ is preserved by $\operatorname{lex}$ it suffices to test the unsatisfiability of $\phi\wedge x\neq y$ for each of these pairs). It follows from the definition of $\operatorname{Feas}$ that for every $g\colon V\to\mathbb{Q}$, if $\phi$ evaluates to a finite value in $\mathfrak{A}$} under the assignment $g$, then $g(x)=g(y)$ for every $(x,y)\in E$. Moreover, for every $(x,y)\in V^{2}\setminus E$, there exists $g\colon V\to\mathbb{Q}$ such that $\phi$ evaluates to a finite value under $g$ and $g(x)\neq g(y)$.
We create a new $\tau$-expression $\phi^{\prime}$ from $\phi$ by replacing each occurrence of $v_{j}$ by $v_{i}$ for every $(v_{i},v_{j})\in E$ such that $i<j$. Let $V^{\prime}$ be the set of variables of $\phi^{\prime}$. By the discussion above, the minimum for $\phi^{\prime}$ over $\mathfrak{A}$} equals the minimum for $\phi$. Moreover, for every $(x,y)\in(V^{\prime})^{2}$, there exists $g^{\prime}\colon V^{\prime}\to\mathbb{Q}$ such that $\phi^{\prime}$ evaluates to a finite value under $g^{\prime}$ and $g^{\prime}(x)\neq g^{\prime}(y)$. Let $\phi^{\prime}:=\phi^{\prime}_{1}+\dots+\phi^{\prime}_{n}$ where for every $j\in[n]$ the summand $\phi^{\prime}_{j}$ is an atomic $\tau$-expression. We execute the following procedure for each $j\in[n]$. Let $\phi^{\prime}_{j}=R(x_{1},\dots,x_{k})$. Let $y_{1}^{j},\dots,y_{\ell_{j}}^{j}$ be an enumeration of all distinct variables that appear in $\{x_{1},\dots,x_{k}\}$ and let $S_{j}$ be a valued relation of arity $\ell_{j}$ defined by $S_{j}(y_{1}^{j},\dots,y_{\ell_{j}}^{j})=R(x_{1},\dots,x_{k})$. Clearly, $S_{j}$ is a minor of $R$. Note that the relation $S_{j}$ might be different for every summand, even if they contain the same relation symbol $R$, due to possibly different variable identifications. Observe that, by the properties of $\phi^{\prime}$, there exists an injective tuple $s^{j}\in\mathbb{Q}^{\ell_{j}}$ such that $S_{j}(s^{j})$ is finite. Note that $S_{j}\in\langle{\mathfrak{A}}\rangle$, and let $m_{j}:=m_{S_{j}}$. By the discussion in the beginning of the proof, $S_{j}(s^{j})=m_{j}$ and $\operatorname{Opt}(S_{j})\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ consists of all tuples that evaluate to $m_{j}$ in $S_{j}$. Since $S_{j}$ attains only finitely many values, we can identify $m_{j}$ in polynomial time for every $j$.
Let $\mathfrak{B}$} be the relational structure with domain $\mathbb{Q}$ and relations $\operatorname{Opt}(S_{1})$, $\dots$, $\operatorname{Opt}(S_{n})$. Let $\psi$ be the instance of $\operatorname{CSP}({\mathfrak{B}})$ obtained from $\phi^{\prime}$ by replacing the summand $\phi^{\prime}_{j}$ by $\operatorname{Opt}(S_{j})(y_{1}^{j},\dots,y_{\ell_{j}}^{j})$ for all $j\in[n]$; all relations in $\psi$ are crisp and hence it can be seen as a primitive positive formula. Note that the variable set of $\psi$ is equal to $V^{\prime}$. By assumption and Lemma 18, $\mathfrak{B}$} is preserved by $\operatorname{ll}$ or $\operatorname{ll}^{*}$. Hence, $\operatorname{CSP}({\mathfrak{B}})$ is in P by Theorem 44. Therefore, the satisfiability of $\psi$ over $\mathfrak{B}$} can be tested in polynomial time. We claim that if $\psi$ is unsatisfiable, then the minimum of $\phi$ is above every rational threshold and the algorithm rejects.
Finally, suppose that there exists a solution $h^{\prime}\colon V^{\prime}\to\mathbb{Q}$ to the instance $\psi$ of $\operatorname{CSP}({\mathfrak{B}})$. Then, for every $j\in[n]$, $\phi^{\prime}_{j}$ takes under $h^{\prime}$ the value $S_{j}(h^{\prime}(y_{1}^{j}),\dots,h^{\prime}(y_{\ell_{j}}^{j}))$. By the definition of $\operatorname{Opt}$, $(h^{\prime}(y_{1}^{j}),\dots,h^{\prime}(y_{\ell_{j}}^{j}))$ minimizes $S_{j}$ and therefore $h^{\prime}$ minimizes $\phi^{\prime}_{j}$. It follows that $h^{\prime}$ minimizes $\phi^{\prime}$ and that the cost of $\phi^{\prime}$ under $h^{\prime}$ is equal to $m_{1}+\dots+m_{n}$. Since the minimum of $\phi^{\prime}$ is equal to the minimum of $\phi$, the algorithm accepts if $m_{1}+\dots+m_{n}\leq u$ and rejects otherwise. This completes the algorithm and its correctness proof. It follows that $\operatorname{VCSP}({\mathfrak{A}})$ is in P. $\qed$
5.3 Expressibility of hard valued relations
In this section we prove several lemmas that are needed to complete the proof of the complexity classification of temporal VCSPs. We introduce the following notation. For $\alpha,\beta,\gamma\in{\mathbb{Q}}\cup\{\infty\}$, define the binary valued relation $R_{\alpha,\beta,\gamma}$ on $\mathbb{Q}$}: